OpenStudy (firejay5):

At liftoff, a 232,000-kg Delta II space launch vehicle is accelerated upward with a force exerted by its rocket engines. This force is called thrust. The Delta II rocket engines produce a thrust of 3.11 million newtons. (a) What is the net force acting on the Delta II at liftoff? (b) What is the acceleration of the Delta II at liftoff? (c) Following liftoff, the acceleration of the Delta II is not constant; it increases. Explain why this happens. (At liftoff, approximately 90% of the weight of the Delta II is rocket fuel.)

OpenStudy (firejay5):

@Harsha19111999

OpenStudy (nincompoop):

c'mon, dude \(\sf F = ma \rightarrow force = mass \times acceleration \)

OpenStudy (harsha19111999):

(a) Net force = Thrust - Gravitational Force = Thrust - mg = 3110000-232000*9.8=836400 N

OpenStudy (firejay5):

I got for a 8.34 X 10^5 N

OpenStudy (nincompoop):

wouldn't that be almost the same?

OpenStudy (harsha19111999):

Okay

OpenStudy (nincompoop):

you have 3 significant figures so follow it

OpenStudy (harsha19111999):

Yeah

OpenStudy (firejay5):

for B, I got 3.60m/s^2

OpenStudy (nincompoop):

do you always want that other people validate your answers before you submit them? that's like cheating

OpenStudy (harsha19111999):

So for (b) Acceleration = Net force / Mass = 3.60 m/sec^2

OpenStudy (firejay5):

exactly that's what I put on my paper

OpenStudy (harsha19111999):

Cool

OpenStudy (harsha19111999):

And (c) Acceleration is not constant. There are many reasons

OpenStudy (firejay5):

I rounded 3.59 to the next whole number

OpenStudy (nincompoop):

you're always asking for validation, I mean every semester

OpenStudy (nincompoop):

I kind of do, because I always see you post questions for the past year

OpenStudy (harsha19111999):

1) The mass of the vehicle may decreased due to the fuel depletion 2) Thrust may increase 3) Gravity variation

OpenStudy (nincompoop):

what if let us say, these posts were to be given to your school for an evaluation, it wouldn't mean anything, yes?

OpenStudy (harsha19111999):

Hey. Cool down everyone. Its not the place to quarrel

OpenStudy (firejay5):

I am sorry @Harsha19111999

OpenStudy (harsha19111999):

No problem. Have confidence in your answers and go ahead. All the best

OpenStudy (firejay5):

thank you! I've actually did research just to figure out these problems

OpenStudy (firejay5):

So for C you would say acceleration isn't constant

OpenStudy (harsha19111999):

Yes

OpenStudy (harsha19111999):

The main reason is the if you travel, fuel burns and comes out. So, the mass of fuel decreases. Acceleration is inversely proportional to Mass. So, acceleration increases